# FIGRET: Fine-Grained Robustness-Enhanced Traffic Engineering

### v1

@@ 4.4. Design DNN structure of FIGRET
The structures primarily used in TE are as follows: Graph Neural Network (GNN) (Xu et al., 2023; Bernárdez et al., 2021), Convolutional Neural Network (CNN) (Valadarsky et al., 2017), and Fully Connected Network (FCN) (Perry et al., 2023). In this section, we briefly explain why we chose FCN over GNN and CNN, with more detail placed in Appendix D.
Network topology handling: no need for GNN. In TE problems, acquiring and utilizing network topology information is crucial for constructing the mapping between TE configurations and MLU. The network topology $G(V,E,c)$, as a graph structure, seems naturally suited for processing with GNN. However, as Function 1 shows, this mapping relationship can be achieved through simple matrix operations, which can be fully handled by FCN. GNN involves processing based on adjacency matrices and node features, which can lead to redundant computations and higher computational complexity, potentially consuming significant memory (Wang et al., 2019; Ying et al., 2018; Zhang et al., 2022). Considering these factors, using FCN as a more efficient alternative seems a more reasonable choice.
CNN vs. FCN: The inappropriateness of CNN. CNN can effectively extract local information in data such as images through convolutional operations. However, in the historical traffic demand data used as input for TE, there is no obvious local information. The traffic sharing a common edge can be spatially close or distant, not necessarily confined to a local area. A detailed example is provided in Figure 11. Therefore, we opted not to use CNN, but rather FNN.
4.5. Coping with link failures
Tunneling protocols are instrumental in identifying paths with failed nodes or links. A traditional approach to addressing link failures in TE involves rerouting traffic around failed paths by allowing traffic sources to proportionally redistribute traffic among their remaining paths (Hong et al., 2013; Jain et al., 2013; Suchara et al., 2011; Perry et al., 2023).
  • In cases where the remaining paths have predefined allocation ratios, the traffic demand from a failed path is proportionally redistributed based on these ratios. Consider a scenario with three paths having allocation ratios of (0.5, 0.3, 0.2). If the first path experiences a failure, the adjusted distribution ratios for the remaining paths would be recalculated to (0, 0.6, 0.4).
  • Conversely, if the remaining paths do not have predefined allocation ratios, the traffic demand from the failed path is equally distributed among them. For instance, in a situation with three paths at ratios (1, 0, 0), and the first path becomes inoperable, the redistributed traffic is allocated evenly, resulting in new ratios of (0, 0.5, 0.5).
@@ 4.5. Coping with link failures
In this section, we evaluate FIGRET. In §5.1, we first describe our evaluation methodology and analyze the types of traffic used in the evaluation to demonstrate that our experiments cover a wide variety of traffic scenarios. Subsequently, in §5.2, we compare FIGRET with state-of-the-art TE schemes, focusing on TE quality, solution times, and precomputation times. Then, the effectiveness of FIGRET in handling link failures and adapting to drifts in traffic patterns is evaluated in §5.3 and §5.4, respectively. Finally, in §5.5, we showcase the adaptive robustness enhancement capabilities of FIGRET, thereby providing an interpretation of FIGRET’s performance.
@@ 5.1. Methodology
Topologies. We consider three WAN topologies, including the pan-European research network, GEANT (Uhlig et al., 2006), as well as two topologies from the Internet Topology Zoo (Knight et al., 2011), namely UsCarrier and Kdl. We also examine various topologies within the data center (DC). This includes the pFabric (Alizadeh et al., 2013), which utilizes a leaf-spine topology, featuring nine Top-of-Rack (TOR) switches. Each ToR switch in the pFabric architecture is connected to all spine switches, forming a network that is logically equivalent to a fully-connected topology. Additionally, we consider two Facebook DC clusters (Roy et al., 2015): the Facebook DB cluster and the Facebook WEB cluster. The Facebook DB cluster comprises MySQL servers that store user data and handle SQL queries, whereas the Facebook WEB cluster is responsible for serving web traffic. For each cluster, we contemplated two types of topologies: the Top-of-Rack (ToR) level and the Point of Delivery (PoD) level. For the PoD level topology, we consider the fat-tree architecture (Al-Fares et al., 2008), whereas for the ToR level topology, the Jellyfish architecture (Singla et al., 2012) is taken into account. Their numbers of nodes and edges are summarized in Table 1. For each topology, we employ Yen’s algorithm to precompute three shortest paths between every pair of nodes as the candidate paths for flow allocation (Abuzaid et al., 2021; Narayanan et al., 2021).
@@
Traffic data. For the GEANT topology, there are publicly available traffic matrix traces, which consist of data aggregated every 15 minutes over four months (Uhlig et al., 2006). In the context of pFabric topology, the pFabric trace is characterized by a Poisson arrival process. When a flow arrives, the source and destination nodes are chosen uniformly at random from a set of 9 different ToR switches. The size of each flow is determined randomly, adhering to the distribution outlined in the ”web search workload” scenario as described in (Alizadeh et al., 2013). In the case of the Facebook topology, there is a public trace of one day’s traffic (Roy et al., 2015). For the PoD-level topology, we aggregate traces into 1-second averaged snapshots of the inter-Pod traffic matrix, while for the ToR-level topology, aggregation is performed into 10-second averaged snapshots of the inter-ToR traffic matrix. The longer aggregation time for ToR-level topologies is attributed to their larger scale, which involves more nodes and a more complex network structure, resulting in greater computational and storage demands for TE. For the UsCarrier and Kdl topologies from Topology Zoo, where no public traffic traces are available, we employ a gravity model (Applegate and Cohen, 2003; Roughan et al., 2002) to generate synthetic traffic.
Traffic characteristics. Our data includes different types of topologies and their corresponding traffic, each with distinct traffic characteristics. By comparing the similarity of the currently-seen TM with historical TMs, we observe whether the traffic is relatively stable and predictable. To verify this property, for every currently-seen TM, we consider a window of historical TMs, find the TMs that most closely resemble this currently-seen TM, and calculate the cosine similarity between the two. Figure 3 shows the distribution of the results. Each candlestick displays the distribution of cosine similarities for different traffics, with the box range extending from the 25th to the 75th percentile. The closer the distribution of a traffic’s cosine similarity is to 1, the more similar it is to past traffic, indicating greater stability. Conversely, a cosine similarity closer to 0 suggests that the traffic is more erratic.
@@
We can identify the following characteristics:
  • The degree of burstiness in traffic increases progressively from WAN traffic to DC POD-level traffic, to DC ToR-level traffic. This may be related to the aggregation of traffic; POD-level traffic aggregates ToR-level traffic, and the WAN connects different data centers, making the aggregation effect more pronounced. The aggregation of multiple traffic streams can cause the bursty variations of different traffic to ’neutralize’ each other, hence the more aggregation there is, the more stable the traffic becomes.
  • Although the WAN traffic generated by the gravity model closely resembles real WAN traffic, the cosine similarity distribution of the traffic generated by the gravity model does not contain anomalies, indicating the traffic is very stable, making TE solutions relatively straightforward.
In summary, the traffic data we have chosen covers a range of traffic characteristics, including stable WAN traffic, highly bursty DC ToR-level traffic, and moderately variable PoD-level traffic. This allows for a thorough testing of the TE solver’s performance in various scenarios.
@@
Baseline. To evaluate the effectiveness of FIGRET improvements relative to other TE schemes, we select the following TE schemes for evaluation: (1) Omniscient TE: This scheme represents an optimization with perfect knowledge of future demands, providing a benchmark for the most efficient performance achievable in all TE schemes. (2) Desensitization-based TE: This is the TE scheme used by Google in Jupiter (Poutievski et al., 2022) and (Teh et al., 2020). This scheme constructs an anticipated matrix composed of the peak values for each source-destination pair within a time window. The optimization of the TE objective is then carried out under the constraint that the sensitivity of each path remains below a certain threshold. (3) Demand-oblivious TE (Applegate and Cohen, 2003): This scheme focuses on optimizing the worst-case performance across all possible traffic demands. (4) Demand-prediction-based TE (Abuzaid et al., 2021; Hong et al., 2013; Jain et al., 2013): This method involves predicting the next incoming traffic demand and configuring accordingly, without considering the mis-predictions that may arise from the traffic uncertainty. (5) COPE (Wang et al., 2006): This scheme enhances demand-oblivious TE by also optimizing over a set of predicted traffic demands. It optimizes MLU across a set of DMs spanned by previously observed DMs while retaining a worst-case performance guarantee. (6) Deep Learning-based TE (DOTE) (Perry et al., 2023): DOTE employs a DNN to directly output the TE configuration based on the traffic demand observed within a given time window. (7) TEAL (Xu et al., 2023): TEAL employs a combination of GNN and Reinforcement Learning (RL) to process a given traffic demand, subsequently outputting a network configuration tailored for this demand. In our experiments, due to the absence of prior knowledge about future traffic, we adopt the network configuration generated based on the traffic demand of the preceding time snap.
@@
Infrastructure and software. Our experiments are carried out on an Intel(R) Xeon(R) Silver 4110 CPU, equipped with 128GB of memory. Furthermore, Nvidia-Tesla P100 GPUs are accessible for all schemes, but only FIGRET, DOTE, and TEAL can utilize them. We implement FIGRET using PyTorch (Paszke et al., 2019). For TE schemes that necessitate solving optimization problems, Gurobi (version 9.5.2) (Gurobi Optimization, 2021) is employed.
@@ 5.2. Comparing with other TE schemes
TE quality. Figure 4 compares the quality of TE solutions between FIGRET and other TE schemes. The y-axis value represents the MLU normalized by that of the Omniscient TE. We note the findings:
@@
  • Desensitization-based TE. Across all topologies, the performance of Desensitization-based TE in terms of both typical MLU (represented by the box) and the worst-case MLU (indicated by the top of the upper dashed line) is not satisfactory. This is due to the fact that Desensitization-based TE imposes unnecessary path sensitivity constraints on non-bursty traffic while providing insufficient path sensitivity constraints for bursty traffic. We will more clearly illustrate this using Figure 6 in §5.5.
@@
  • Demand-prediction-based TE. Demand-prediction-based Traffic TE methods primarily excel in environments with stable traffic demands. However, their effectiveness significantly diminishes in sudden or unexpected traffic bursts. The performance of these methods is highly dependent on the accuracy of traffic demand predictions.
@@
  • Demand-oblivious TE & COPE. Due to the extensive computation time and memory requirements of Demand-oblivious TE and COPE, we are only able to use a few smaller-scale topologies. The two methods for handling mis-prediction are essentially the use of a set-based approach as introduced in §4.1. It is evident that selecting a set not only increases computational complexity but also makes it difficult to achieve a good balance between typical performance and robustness to traffic bursts.
@@
  • DOTE. DOTE, the best-performing algorithm in existing TE (Perry et al., 2023), leverages deep learning to link historical traffic data with imminent network configurations, achieving outstanding typical performance. DOTE implicitly learns to achieve the lowest expected MLU with $D^{\text{expect}}$, reaching the optimal average performance among existing methods. However, its drawback lies in handling burst situations. For example, in the GEANT data, as shown in Figure 3, while most of the data points have high cosine similarity, close to 1, there are some outliers where the traffic’s cosine similarity with past window traffic is very low, representing exceptional cases. This leads to DOTE’s MLU performance on GEANT data being close to that of Omniscient in most situations but with some high values. As illustrated in Figure 4(b), in ToR-level DC with high dynamic characteristics, DOTE performs worse than FiGRET in terms of both average performance and robustness.
@@
  • TEAL. TEAL is designed to train a fast and scalable TE scheme. It receives a traffic demand and outputs a network configuration specifically for this demand. However, when this network configuration is applied to future traffic demands, it tends to underperform.
@@
  • FIGRET. Our designed FIGRET achieves better results than previous TE methods, striking a better balance between average performance and robustness. For instance, compared to the TE scheme currently in production by Google’s Jupiter, FIGRET reduces the average MLU by 9%-34%. Against the state-of-the-art DOTE algorithm, specifically designed to lower MLU, FIGRET achieves a reduction in average MLU of 4.5% and 5.3% on ToR-level Facebook DB and ToR-level Facebook WEB, respectively, both characterized by high traffic dynamics. Even in stable topologies like Pod-level Datacenters, FIGRET does not underperform compared to DOTE. Simultaneously, we consider situations where the normalized MLU is greater than 2 as instances of severe congestion caused by inadequate network configuration. For the ToR-level Facebook DB, FIGRET exhibits a 41% lower incidence of severe congestion compared to DOTE, while for the ToR-level Facebook WEB, the reduction is 53.9%.
(a) GEANT &\&& pFabric
(b) ToR-level Facebook DB &\&& ToR-level Facebook WEB
(c) PoD-level Facebook DB &\&& PoD-level Facebook WEB
(d) Cogentco &\&& UsCarrier
Figure 4. Performance of FIGRET and baselines under the objective of minimizing max link utilization (MLU).
Solver time. We consider the solver time to consist of two parts: the calculation time required for applying each TE scheme to a new demand matrix, and the required precomputation time. Table 2 presents a comparison of the calculation time for different TE schemes. Table 2 lists four items: FIGRET, a Linear Programming (LP) approach that optimizes solely for expected traffic demands without applying sensitivity constraints, Desensitization-based TE, which addresses the optimization problem by imposing constraints to limit path sensitivity, and Oblivious & COPE, which are precomputed TE solutions not updated thereafter; hence, we only denote whether they can solve for the corresponding network. It is evident that a trained neural network, using matrix operations, can quickly compute a TE configuration. The comparison between LP and Des TE shows that adding constraints to the optimization problem to ensure robustness increases the solution time. However, by incorporating a penalty term into the neural network, we enhance robustness without increasing the latency. Compared to Des TE, FIGRET achieves a speed-up of 35×35$\times$35 × to 1800×1800$\times$1800 ×.
Network Calculation time (s)
FIGRET LP Des TE OBL&Cope
GEANT (#Nodes 23) 0.002 0.04 0.07 Feasible
ToR DB (#Nodes 155) 0.005 1.60 5.00 Infeasible
ToR WEB (#Nodes 324) 0.009 7.30 17.00 Infeasible
@@
Table 2. Comparing the calculation time across various TE schemes. OBL represents Oblivious.
@@
Regarding the training time, since we employ a simple FCN architecture in FIGRET, the duration remains within an acceptable range. COPE and Demand-oblivious require substantial precomputation latency and a significant amount of memory. Given a maximum runtime of 1 day and machine memory constraints, they only completed the topologies for GEANT, pFabric, and Pod-level DC. Past public results also show their applicability is limited to smaller topologies (Wang et al., 2006; Perry et al., 2023). Simultaneously, compared to TEAL, which employs reinforcement learning and graph neural networks, FIGRET also has advantages in terms of training time. Detailed data are presented in Table 8.

### v3

@@ 4.4. Design DNN structure of FIGRET
The structures primarily used in TE are as follows: Graph Neural Network (GNN) (Xu et al., 2023; Bernárdez et al., 2021), Convolutional Neural Network (CNN) (Valadarsky et al., 2017), and Fully Connected Network (FCN) (Perry et al., 2023). In this section, we briefly explain why we chose FCN over GNN and CNN, with more detail available in Appendix D.
Network topology handling: no need for GNN. In TE problems, acquiring and utilizing network topology information is crucial for constructing the mapping between TE configurations and MLU. The network topology $G(V,E,c)$, as a graph structure, seems naturally suited for processing with GNN. However, as Function 1 in Appendix D.1 shows, this mapping relationship can be achieved through simple matrix operations, which can be fully handled by FCN. GNN involves processing based on adjacency matrices and node features, which can lead to redundant computations and higher computational complexity, potentially consuming significant memory (Wang et al., 2019; Ying et al., 2018; Zhang et al., 2022). Considering these factors, using FCN seems a more reasonable choice.
CNN vs. FCN: The inappropriateness of CNN. CNN can effectively extract local information in data such as images through convolutional operations. However, in the historical traffic demand data used as input for TE, there is no obvious local information. The traffic sharing a common edge can be spatially close or distant, not necessarily confined to a local area. A detailed example is provided in Figure LABEL:fig:_cnn. Therefore, we opted not to use CNN, but rather FCN.
4.5. Coping with node/link failures
In large networks, node/link failures are inevitable, causing some paths unavailable to serve traffic. A widely-adopted approach to addressing network failures in TE involves rerouting traffic around failed paths by allowing traffic sources to proportionally redistribute traffic among their remaining paths (Hong et al., 2013; Jain et al., 2013; Suchara et al., 2011; Perry et al., 2023).
  • In cases where the remaining paths have allocation ratios, the traffic demand from a failed path is proportionally redistributed based on these ratios. Consider a scenario with three paths having allocation ratios of (0.5, 0.3, 0.2). If the first path experiences a failure, the adjusted distribution ratios for the remaining paths would be set as (0, 0.6, 0.4).
  • Conversely, if the remaining paths do not have allocation ratios, the traffic demand from the failed path is equally distributed among them. For instance, in a situation with three paths at ratios (1, 0, 0), and the first path fails, the redistributed traffic is allocated evenly, resulting in new ratios of (0, 0.5, 0.5).
@@ 4.5. Coping with node/link failures
In this section, we evaluate FIGRET. In §5.1, we first describe our evaluation methodology and analyze the types of traffic used in the evaluation to demonstrate that our experiments cover a wide variety of traffic scenarios. Subsequently, in §5.2, we compare FIGRET with state-of-the-art TE schemes, focusing on TE quality, solution times, and precomputation times. Then, the effectiveness of FIGRET in handling link failures and adapting to drifts in traffic patterns is evaluated in §5.3 and §5.4, respectively. Finally, in §5.5, we showcase the fine-grained robustness enhancement capabilities of FIGRET, thereby providing an interpretation of FIGRET’s performance.
@@ 5.1. Methodology
Topologies. We consider three WAN topologies, including the pan-European research network, GEANT (Uhlig et al., 2006), as well as two topologies from the Internet Topology Zoo (Knight et al., 2011), namely UsCarrier and Kdl. We also examine various data center topologies. This includes the pFabric (Alizadeh et al., 2013), which utilizes a leaf-spine topology with nine Top-of-Rack (TOR) switches. We change the topology from leaf-spine to a fully connected direct-connect network, because TE is rarely used in leaf-spine topologies. Additionally, we consider two Meta’s DC clusters (Roy et al., 2015): the Meta DB cluster and the Meta WEB cluster. The Meta DB cluster comprises MySQL servers that store user data and handle SQL queries, whereas the Meta WEB cluster is responsible for serving web traffic. For each cluster, we contemplate two types of topologies: the Top-of-Rack (ToR) level and the Point of Delivery (PoD) level. For Meta’s data centers, we also adopt the direct-connect network as the topology. For the PoD level topology, we consider the fully connected network, whereas, for the ToR level topology, we utilize the random regular graph (Singla et al., 2012). Their numbers of nodes and edges are summarized in Table 1. For each topology, unless otherwise stated, we employ Yen’s algorithm to precompute the three shortest paths between every pair of nodes as the candidate paths for flow allocation (Abuzaid et al., 2021; Narayanan et al., 2021).
@@
Traffic data. For the GEANT topology, there are publicly available traffic matrix traces, which consist of data aggregated every 15 minutes over four months (Uhlig et al., 2006). In the context of pFabric topology, the pFabric trace is characterized by a Poisson arrival process. When a flow arrives, the source and destination nodes are chosen uniformly at random from the different ToR switches. The size of each flow is determined randomly, adhering to the distribution outlined in the “web search workload” scenario as described in (Alizadeh et al., 2013). In the case of the Meta topology, there is a public trace of one day’s traffic (Roy et al., 2015). For the PoD-level topology, we aggregate traces into 1-second snapshots of the inter-Pod traffic matrix, while for the ToR-level topology, aggregation is performed into 10-second snapshots of the inter-ToR traffic matrix. The longer aggregation time for ToR-level topologies is attributed to their larger scale, which involves more nodes and a more complex network structure, resulting in greater computational and storage demands for TE. For the UsCarrier and Cogentco topologies from Topology Zoo, where no public traffic traces are available, we employ a gravity model (Applegate and Cohen, 2003; Roughan et al., 2002) to generate synthetic traffic.
Traffic characteristics. Our data includes different types of topologies and their corresponding traffic, each with distinct traffic characteristics. By comparing the similarity of the currently-seen TM with historical TMs, we observe whether the traffic is relatively stable and predictable. Specifically, for every currently-seen TM, we consider a window of historical TMs, find the TMs that most closely resemble this currently-seen TM, and calculate the cosine similarity between the two. Figure 9 shows the distribution of the results. Each candlestick displays the distribution of cosine similarities for different traffics, with the box range extending from the 25th to the 75th percentile. The closer the distribution of a traffic’s cosine similarity is to 1, the more similar it is to past traffic, indicating greater stability. Conversely, a cosine similarity closer to 0 suggests that the traffic is more erratic. We can identify the following characteristics:
  • The degree of burstiness in traffic increases progressively from WAN traffic to DC POD-level traffic, to DC ToR-level traffic. This may be related to the aggregation of traffic; POD-level traffic aggregates ToR-level traffic, and the WAN connects different data centers, making the aggregation effect more pronounced. The aggregation of multiple traffic streams can cause the bursty variations of different traffic to ‘neutralize’ each other, hence the more aggregation there is, the more stable the traffic becomes.
  • The WAN traffic generated by the gravity model resembles the trend of real WAN traffic. However, it cannot capture all the characteristics of real traffic, as it fails to reflect traffic bursts. In our experiments, we use it to understand the TE performance in stable traffic conditions.
In summary, the traffic data exhibits various characteristics, including profiles with low, moderate, and high burstiness. This allows for a thorough evaluation of the TE performance in various scenarios.
@@
Baseline. We select the following TE schemes to compare with FIGRET: (1) Omniscient TE: This scheme represents an optimization with perfect knowledge of future demands, providing a benchmark for the most efficient performance achievable in all TE schemes. (2) Desensitization-based TE: This is the TE scheme used by Google in Jupiter data center(Poutievski et al., 2022) and (Teh et al., 2020). This scheme constructs an anticipated matrix composed of the peak values for each source-destination pair within a time window. Then it optimizes the TE objective under the constraint that path sensitivity remains below a predetermined threshold. (3) Demand-oblivious TE (Applegate and Cohen, 2003): This scheme focuses on optimizing the worst-case performance across all possible traffic demands. (4) Demand-prediction-based TE (Abuzaid et al., 2021; Hong et al., 2013; Jain et al., 2013): This method involves predicting the next incoming traffic demand and configuring accordingly, without considering the mis-predictions that may arise from the traffic uncertainty. (5) COPE (Wang et al., 2006): This scheme enhances demand-oblivious TE by also optimizing over a set of predicted traffic demands. It optimizes MLU across a set of DMs predicted based on previously observed DMs while retaining a worst-case performance guarantee. (6) Deep Learning-based TE (DOTE) (Perry et al., 2023): DOTE employs a DNN to directly output the TE configuration based on the traffic demand observed within a given time window. (7) TEAL (Xu et al., 2023): TEAL employs a combination of GNN and Reinforcement Learning (RL) to process a given traffic demand, subsequently outputting a network configuration tailored for this demand. In our experiments, due to the absence of prior knowledge about future traffic, we apply the TE solution computed from the traffic demand of the preceding time snapshot to the next time snapshot. (8) SMORE (Kumar et al., 2018): SMORE employs Räcke’s oblivious routing paths (Racke, 2002) for path selection, with traffic splitting ratios optimized for the predicted future traffic demands.
@@
Infrastructure and software. Our experiments are carried out on an Intel(R) Xeon(R) Silver 4110 CPU, equipped with 128GB of memory. Furthermore, Nvidia-Tesla P100 GPUs are accessible for all schemes, but only FIGRET, DOTE, and TEAL can utilize them. We implement FIGRET using PyTorch (Paszke et al., 2019). For detailed information on neural network architecture, parameters, and optimizer settings, please refer to Appendix D. For TE schemes that necessitate solving optimization problems, Gurobi (version 9.5.2) (Gurobi Optimization, 2021) is employed.
@@ 5.2. Comparing with other TE schemes
TE quality. Figure LABEL:fig:_performance_of_figret compares the quality of TE solutions between FIGRET and other TE schemes. The y-axis value represents the MLU normalized by that of the Omniscient TE. We note the findings:
@@
  • Desensitization-based TE. Across all topologies, the performance of Desensitization-based TE in terms of both normal-case MLU (represented by the box) and the worst-case MLU (indicated by the top of the upper dashed line) is not satisfactory. This is due to the fact that Desensitization-based TE imposes unnecessary path sensitivity constraints on non-bursty traffic while providing insufficient path sensitivity constraints for bursty traffic. We will more clearly illustrate this using Figure LABEL:fig:_interpret_termol in §5.5.
@@
  • Demand-prediction-based TE. Demand-prediction-based TE methods primarily excel in environments with stable traffic demands. However, their effectiveness significantly diminishes when unexpected traffic bursts occur. The performance of these methods is highly dependent on the accuracy of traffic demand predictions.
@@
  • Demand-oblivious TE & COPE. Due to the extensive computation time and memory requirements of Demand-oblivious TE and COPE, we can only evaluate a few smaller-scale topologies. The two methods for handling mis-prediction are essentially the use of a set-based approach as introduced in §4.1. It is evident that selecting a set not only increases computational complexity but also makes it difficult to achieve a good balance between normal-case performance and robustness to traffic bursts.
@@
  • DOTE. DOTE, the best-performing algorithm in existing TE (Perry et al., 2023), leverages deep learning to map historical traffic data with the next network configurations, achieving outstanding normal-case performance. DOTE implicitly learns to achieve the lowest expected MLU with $D^{\text{expect}}$, reaching the optimal average performance among existing methods. However, its drawback lies in handling burst situations. For example, in the GEANT data, as shown in Figure 9, while most of the data points have high cosine similarity, close to 1, there are some outliers where the traffic’s cosine similarity with past window traffic is very low, representing the unexpected burst scenarios cases. This leads to DOTE’s MLU performance on GEANT data being close to that of Omniscient in most situations but with some peak values. As illustrated in Figure LABEL:fig:_tor_a_tor_b, in ToR-level DC with high dynamic characteristics, DOTE performs worse than FiGRET in terms of both average performance and robustness.
@@
  • TEAL. TEAL is designed to train a fast and scalable TE scheme. It receives a traffic demand and outputs a network configuration specifically for this demand. However, when this network configuration is applied to traffic demands with unexpected bursts, it tends to underperform.
@@
  • FIGRET. Our designed FIGRET achieves better results than previous TE methods, striking a better balance between normal-case performance and burst-case performance. For instance, compared to the TE scheme currently in Google’s production data center network, FIGRET reduces the average MLU by 9%-34%. Against the state-of-the-art DOTE algorithm, FIGRET achieves a reduction in average MLU of 4.5% and 5.3% on ToR-level Meta DB and ToR-level Meta WEB, respectively, both characterized by high traffic dynamics. Even in stable topologies like WAN or Pod-level Datacenters, FIGRET performs no worse than DOTE. Simultaneously, we consider situations where the normalized MLU is greater than 2 as instances of severe congestion caused by inadequate network configuration. For the ToR-level Meta DB, FIGRET exhibits a 41% lower incidence of severe congestion compared to DOTE, while for the ToR-level Meta WEB, the reduction is 53.9%.
Furthermore, as shown in Figure LABEL:fig:_zoo, all TE methods exhibit a low normalized MLU with no peaks. This is because the synthetic traffic generated by the gravity model is very stable and lacks traffic bursts. That said, handling traffic bursts is still critical for real WAN traffic because real WAN traffic does contain unexpected bursts.
Network Calculation time (s) Precomp. time (s)
FIGRET LP Des TE Oblivious & Cope FIGRET TEAL Oblivious COPE
GEANT (#Nodes 23, #Edges 74) 0.002 0.04 0.07 Feasible 150 1500 100 1000
ToR DB (#Node 155, #Edges 7194) 0.005 1.60 5.00 Infeasible 500 7000 - -
ToR WEB (#Node 324, #Edges 31520) 0.009 7.30 17.00 Infeasible 1500 20000 - -
@@
Table 2. Comparing the calculation and precomputation time across various TE schemes.
Solver time. The solver time consists of two parts: the calculation time required to compute a new TE scheme for each new demand matrix, and the precomputation time. Table 2 presents a comparison of the calculation time for different TE schemes. Table 2 lists four items: 1) FIGRET, 2) a Linear Programming (LP) approach that optimizes solely for expected traffic demands without applying any anti-burst strategy, 3) Desensitization-based TE, which handles traffic bursts by limiting path sensitivity, and 4) Oblivious & COPE, which precompute TE solutions but do not update them thereafter; hence, we only test whether they can solve a problem instance. Clearly, FIGRET has the smallest calculation time, while Oblivious & COPE TE schemes have the poorest scalability, failing to compute a solution for larger networks. The comparison between LP and Des TE shows that adding the path sensitivity constraints to handle traffic bursts increases the LP solver calculation time. In contrast, thanks to the adoption of deep neural network, FIGRET is able to enhance robustness without increasing solver complexity. Compared to Des TE, FIGRET achieves a speed-up of 35×35$\times$35 × to 1800×1800$\times$1800 ×.
@@
Regarding the training time, since we employ a simple FCN architecture in FIGRET, the precomputation time required is shorter than COPE and reinforcement learning-based TE, especially when dealing with large-scale topologies. COPE and Demand-oblivious require substantial precomputation latency and a significant amount of memory. Given a maximum runtime of 1 day and machine memory constraints, they only completed the topologies for GEANT, pFabric, and Pod-level DC. Past public results also show their applicability is limited to smaller topologies (Wang et al., 2006; Perry et al., 2023). Simultaneously, compared to TEAL, which employs reinforcement learning and graph neural networks, FIGRET also has advantages in terms of training time. Detailed data are presented in Table 2.
Path selection. We compare FIGRET with SMORE. SMORE uses a set of paths computed by Räcke’s oblivious routing algorithm. To ensure fairness, we use the same Räcke’s routing algorithm to generate paths for other algorithms. In this case, the prediction-based TE coincides with SMORE. The results are summarized in Figure 13. Again, FIGRET demonstrates the best performance among all the TE algorithms. By comparing Figure 13 and Figure LABEL:fig:_GEANT_pFabric, it is evident that path selections do not fundamentally affect the effectiveness of different TE schemes. SMORE aims to enhance the robustness merely through path selection. After the paths have been determined, SMORE optimizes the TE performance based solely on the predicted traffic, without considering the potential traffic bursts. As shown by the Pred TE results in the GEANT network in Figure 13, SMORE incurs the highest tail normalized MLU among all the TE schemes. This indicates that merely enhancing TE robustness through path selection is insufficient.
Figure 13. Performance of FIGRET and baselines with routing paths chosen by SMORE. “Pred TE” represents SMORE.